OpenStudy (anonymous):

What's the difference between "Linear Combination," and the "Substitution Method," and how do I solve using Substitution?

OpenStudy (anonymous):

for substitution you express one variable through another. For combination (elimination) you eliminate one variable.

OpenStudy (anonymous):

To solve using the substitution method, we rewrite one equation and sub it into the other. For example 4x + 3y = (-2) [1] 3x + 2y = (-3) [2] Let rewrite [1] so y is our subject 4x+3y=-2 3y=-2-4x y=-2/3-(4/3)x Now that we have y sub this into the second equation and solve for x 3x+2(-2/3-(4/3)x)=-3 3x-4/3-8/3 x=-3 1/3 x=-5/3 x=-5 Now sub x=-5 into our rearranged y equation y=-2/3-(4/3)*(-5) =-2/3+20/3 =6 So y=6 and x=-5